# RaFE.github.io

### v1

@@ Abstract
NeRF (Neural Radiance Fields) has demonstrated tremendous potential in novel view synthesis and 3D reconstruction, but its performance is sensitive to input image quality, which struggles to achieve high-fidelity rendering when provided with low-quality sparse input viewpoints. Previous methods for NeRF restoration are tailored for specific degradation type, ignoring the generality of restoration. To overcome this limitation, we propose a generic radiance fields restoration pipeline, named RaFE, which applies to various types of degradations, such as low resolution, blurriness, noise, compression artifacts, or their combinations. Our approach leverages the success of off-the-shelf 2D restoration methods to recover the multi-view images individually. Instead of reconstructing a blurred NeRF by averaging inconsistencies, we introduce a novel approach using Generative Adversarial Networks (GANs) for NeRF generation to better accommodate the geometric and appearance inconsistencies present in the multi-view images. Specifically, we adopt a two-level tri-plane architecture, where the coarse level remains fixed to represent the low-quality NeRF, and a fine-level residual tri-plane to be added to the coarse level is modeled as a distribution with GAN to capture potential variations in restoration. We validate RaFE on both synthetic and real cases for various restoration tasks, demonstrating superior performance in both quantitative and qualitative evaluations, surpassing other 3D restoration methods specific to single task. Please see our project website RaFE.github.io.
@@ 4.2 Results
Quantitative Results. We conduct extensive quantitative comparisons with various baselines across different restoration task in Tab. 1(a) for super-resolution, Tab. 2(b) and Tab. 2(a) for deblurring, Tab. 2(c) for denoising and Tab. 1(b) for mixed degradation. As analyzed before, although most of the time our method falls slightly behind on the reconstruction metrics like PSNR and SSIM compared with other baselines, which only measure local pixel-aligned similarity between the rendered novel views and the ground truth images, are less indicative since uncertainties naturally exist in generative procedure. Taking the super-resolution results on Blender data as an example (Tab. 1(a)), the simplest baseline NeRF-Perframe have already achieved the best reconstruction metrics, but as shown in Figure. 3(a), its visual quality is vastly inferior compared with our results. Through the error map, we found the mis-alignment between the generated 3D and input 3D causes the drop of PSNR and SSIM. By contrast, on the perceptual metrics like LPIPS and non-reference based metrics including LIPE and MANIQA, which could more effectively reflect the restoration performance, our method consistently achieves better results when compared with other baselines, demonstrating its clear advantages.
@@
Geometry Refinement. Compared with previous NeRF restoration methods, most of which focusing on using 2D refinement to resolve frame-wise defects, our method firstly leverages the general priors of large foundation models and GAN, achieving the open-domain 3D-based restoration. Thus, our restoration algorithm could not only recover the high-quality and view-consistent images, but also refine the 3D geometry. We demonstrate this advantage using the Ficus data of Blender, as shown in Fig. 5(a), through comparing the extracted mesh from the restored NeRF models, our method effectively learns the restoration in the 3D tri-plane space and recover better geometry.
@@ 4.4 Discussion
Effects of different restoration models. To investigate the influence of different 2D restoration models on our method, we tested two additional off-the-shelf restoration models for the super-resolution task, including diffusion-based DiffBIR [25] and non-diffusion-based SRFormer [53]. As shown in Fig. 5(b), diffusion-based model DiffBIR shows larger restoration diversity over SRFormer as expected by measuring the diversity score. When the repaired images exhibit diversity, direct reconstruction inevitably leads to the blurriness to varying degrees due to the exisiting multi-view inconsistency. By contrast, through modeling the distribution of the potential high-quality NeRFs, our method successfully accommodates these inconsistencies and consistently achieves better performance over NeRF-Perframe, demonstrating the great generalization capability of RaFE to different 2D restoration model.

### v2

# zkaiwu.github.io/RaFE

@@ Abstract
NeRF (Neural Radiance Fields) has demonstrated tremendous potential in novel view synthesis and 3D reconstruction, but its performance is sensitive to input image quality, which struggles to achieve high-fidelity rendering when provided with low-quality sparse input viewpoints. Previous methods for NeRF restoration are tailored for specific degradation type, ignoring the generality of restoration. To overcome this limitation, we propose a generic radiance fields restoration pipeline, named RaFE, which applies to various types of degradations, such as low resolution, blurriness, noise, compression artifacts, or their combinations. Our approach leverages the success of off-the-shelf 2D restoration methods to recover the multi-view images individually. Instead of reconstructing a blurred NeRF by averaging inconsistencies, we introduce a novel approach using Generative Adversarial Networks (GANs) for NeRF generation to better accommodate the geometric and appearance inconsistencies present in the multi-view images. Specifically, we adopt a two-level tri-plane architecture, where the coarse level remains fixed to represent the low-quality NeRF, and a fine-level residual tri-plane to be added to the coarse level is modeled as a distribution with GAN to capture potential variations in restoration. We validate RaFE on both synthetic and real cases for various restoration tasks, demonstrating superior performance in both quantitative and qualitative evaluations, surpassing other 3D restoration methods specific to single task. Please see our project website zkaiwu.github.io/RaFE.
@@ 4.2 Results
Quantitative Results. We conduct extensive quantitative comparisons with various baselines across different restoration task in Tab. 0(a) for super-resolution, Tab. 2 and Tab. 2 for deblurring, Tab. 2 for denoising and Tab. 0(b) for mixed degradation. As analyzed before, although most of the time our method falls slightly behind on the reconstruction metrics like PSNR and SSIM compared with other baselines, which only measure local pixel-aligned similarity between the rendered novel views and the ground truth images, are less indicative since uncertainties naturally exist in generative procedure. Taking the super-resolution results on Blender data as an example (Tab. 0(a)), the simplest baseline NeRF-Perframe have already achieved the best reconstruction metrics, but as shown in Figure. 2(a), its visual quality is vastly inferior compared with our results. Through the error map, we found the mis-alignment between the generated 3D and input 3D causes the drop of PSNR and SSIM. By contrast, on the perceptual metrics like LPIPS and non-reference based metrics including LIPE and MANIQA, which could more effectively reflect the restoration performance, our method consistently achieves better results when compared with other baselines, demonstrating its clear advantages.
@@
Geometry Refinement. Compared with previous NeRF restoration methods, most of which focusing on using 2D refinement to resolve frame-wise defects, our method firstly leverages the general priors of large foundation models and GAN, achieving the open-domain 3D-based restoration. Thus, our restoration algorithm could not only recover the high-quality and view-consistent images, but also refine the 3D geometry. We demonstrate this advantage using the Ficus data of Blender, as shown in Fig. 5, through comparing the extracted mesh from the restored NeRF models, our method effectively learns the restoration in the 3D tri-plane space and recover better geometry.
[…]
@@ 4.4 Discussion
Effects of different restoration models. To investigate the influence of different 2D restoration models on our method, we tested two additional off-the-shelf restoration models for the super-resolution task, including diffusion-based DiffBIR [25] and non-diffusion-based SRFormer [53]. As shown in Fig. 5, diffusion-based model DiffBIR shows larger restoration diversity over SRFormer as expected by measuring the diversity score. When the repaired images exhibit diversity, direct reconstruction inevitably leads to the blurriness to varying degrees due to the exisiting multi-view inconsistency. By contrast, through modeling the distribution of the potential high-quality NeRFs, our method successfully accommodates these inconsistencies and consistently achieves better performance over NeRF-Perframe, demonstrating the great generalization capability of RaFE to different 2D restoration model.